4. Suppose a firm produces according to the production function Q = AL0.6K0.2, and faces wage rate ₵20, a rental cost of capital ₵10, and sells output at a price of ₵40.

a. Obtain and expression for the factor demand functions if A=2.

b. Compute the profit-maximizing factor demands for capital and labour if A=2.

5. In the short run, a competitive firm has a production function,

Q = f(L) = 2.6667L0.75. The output price is $8 per unit and the wage is $5 per hour. Find the short-

run labor demand curve of the firm.



1
Expert's answer
2021-11-01T17:55:29-0400

Solution:

a.). Q = AL0.6K0.2

A = 2

Q = 2L0.6K0.2

Input demand for labor, L = F(w,r,p)

Input demand for Capital, K = F(w,r,p)

 

b.). Profit (Ï€) = TR = TC

TR = P "\times" Q = 40(2L0.6K0.2)

TC = wL + rK = 20L + 40K

Profit (π) = 40(2L0.6K0.2) – 20L – 40K

Profit (π) = 80L0.6K0.2) – 20L – 40K

"\frac{\partial \pi } {\partial L}" = 48L-0.4K0.2 – 20 = 0                   (a)


"\frac{\partial \pi } {\partial K}" =16L0.6K-0.8 – 40 = 0                   (b)


Solve for equation (a) for L:

48L-0.4K0.2 = 20

"\frac{48K } { L^{0.4} }"= 20      

48K = 20L0.4

2.4K = L0.4

L = 8.92K2.5

Plug into equation (b):

16L0.6K-0.8 = 40

16(8.92K2.5)0.6K-0.8 = 40

142.72K1.5K-0.8 = 40

"\frac{142.72K^{1.5} } { K^{0.8} }" = 40

142.72K1.5 = 40K0.8

K = 0.16

L = 8.92K2.5 = 8.92(0.162.5) = 0.09

Profit maximizing factor demands for labor and capital = (0.09, 0.16)


c.). Q = f(L) = 2.6667L0.75

MRPL = MR "\times" MPL = w

MR = P = 8

W = 5

MPL = "\frac{\partial Q} {\partial L}"= 2.000025L-0.25

MRPL = 8 "\times" 2.000025L-0.25 = 16.0002L-0.25

16.0002L-0.25 = w

16.0002L-0.25 = 5


"\frac{16.0002}{L^{0.25} }" = 5


16.0002 = 5L0.25


L = 104.86


The short-run labor demand curve of the firm = 104.86